# Machine Unlearning Through Fine-Grained Model Parameters Perturbation

### v3

@@ Abstract
However, commonly used training data are independent and identically distributed, for inexact machine unlearning, current metrics are inadequate in quantifying forgetting degree that occurs after unlearning. To address this quantification issue, we introduce SPD-GAN, which subtly perturbs data distribution targeted for unlearning. Then, we evaluate unlearning degree by measuring the performance difference of the models on the perturbed unlearning data before and after unlearning. Furthermore, to demonstrate efficacy, we tackle the challenge of evaluating machine unlearning by assessing model generalization across unlearning and remaining data. To better assess the unlearning effect and model generalization, we propose novel metrics, namely, the forgetting rate and memory retention rate. By implementing these innovative techniques and metrics, we achieve computationally efficacious privacy protection in machine learning applications without significant sacrifice of model performance. A by-product of our work is a novel method for evaluating and quantifying unlearning degree.
@@ I Introduction
Users may want the removal of personal data in possession of an organization, typically deleting it randomly based on their specific needs. In several jurisdictions, this has compliance implications under regulations such as the General Data Protection Regulation (GDPR) in the European Union [1] and the California Consumer Privacy Act in the United States [2]. Machine learning models trained using such data retain their influence, thus risking privacy compromises [3]. For instance, in a recommendation system using collaborative filtering [4], if the model does not eliminate the influence of removed data (which we call ‘unlearning data’), it may continue to make recommendations based on user similarity.
@@
Differently, inexact machine unlearning [10, 11] aims to reduce time and resources requirements by selectively eliminating the data’s influence on the model. Current inexact machine unlearning methods based on model weight perturbation typically involve adding random Gaussian noise to all model parameters. We propose to achieve inexact unlearning using a more fine-gained perturbation. To that end, we design two inexact machine unlearning strategies: Random-k and Top-K. By selectively perturbing a small subset of parameters, we aim to achieve the desired effect of machine unlearning efficiently while also reducing the impact on the overall model performance.
@@
However, the existing metrics are unable to reveal the extent of unlearning. Suppose splitting training data into two parts: unlearning data $D_{UL}$ and remaining data $D_{RE}$ $\left(D=D_{UL}\cup D_{RE},len(D_{UL})\ll len(D_{RE})\right)$. As shown in Fig.1, which shows an early glimpse of some results achieved by one of our own unlearning techniques), while the accuracy may vary across $D_{UL}$ and $D_{RE}$ for each model, it consistently maintains a high accuracy for $D_{UL}$. Consequently, it becomes challenging to quantify the extent to which the model eliminates the influence of $D_{UL}$. Since the performance difference of unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$ is relatively small. We attribute this to two reasons, first, a good generalization capability of the model, second, the characteristic that the training data followed an independent and identically distributed (i.i.d) pattern.
In order to amplify the performance difference between the unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$, we propose a SPD-GAN (Slightly Perturbed Distribution-Generative Adversarial Network) to break the i.i.d attribute of the unlearning data. For the unlearning data $D_{UL}$, we introduce perturbations to it’s distribution, thus we gain perturbed unlearning data $D_{p}$. We limit the perturbations to be very slight so that the accuracy of source model $M$ on $D_{p}$ is close to the accuracy of $M$ on $D_{UL}$. Suppose $P(M,D)$ represents the performance of model $M$ on data $D$, seeing in the Equation.1, if the perturbed unlearning data $D_{p}$ exhibits significant performance difference between source model and unlearning model, we can then obtain the unlearning degree of each unlearning strategies.
@@ II-A User Privacy
Federated learning (FL) is a decenterlized approach to protect user privacy, where training data is distributed among multiple devices and servers, and the model is trained in a decentralized manner without centralizing the data [26]. FL methods based on differential privacy have been proposed to protect user privacy, with the principal focus on attacks and defenses [27] in FL systems. Furthermore, recent advancements in FL incorporate various cryptographic techniques, such as secure multi-party computation and homomorphic encryption, to enhance the security and privacy of the training process. This ongoing research strengthens the privacy-preserving capabilities of FL.
@@ II-B Machine Unlearning
Several approaches have emerged to address the limitations of retraining models from scratch for machine unlearning, since it is generally impractical. The most typical approach is SISA (Sharded, Isolated, Sliced, and Aggregated) training [6], which splits data into shards and trains them in isolation for speed and effectiveness when unlearning. Liu et al. [29] combine continual learning with private learning based on the SISA idea, but SISA-based methods often consume a lot of storage. Following the combination of machine unlearing and continual learning, Zuo et al. [30] propose a embedding-based framework, which modifies the data rather than the model itself to achieve unlearning goal by using vector databases.
@@
In addition to combining with continual learning, machine unlearning is also integrated with other approaches, such as federated learning, forming a research area known as federated unlearning [32] [33]. In federated unlearning, considerations are made for clients leaving the federation or for unlearning data corresponding to clients. For instance, Liu et al. [32] propose FedEraser, which unlearns a client’s data by reconstructing an unlearned global model instead of retraining from scratch. Wu et al. [33] utilize knowledge distillation techniques to recover the contribution of client models to achieve federated unlearning.
@@ II-C Adversarial Learning
Moreover, these techniques have laid the groundwork for numerous subsequent studies in adversarial machine learning, illustrating the continuous evolution of this field. In other words, leveraging adversarial training enhances the robustness of the model and better identify and defend against malicious attacks. Smoothed ViTs [42] are trained using only one columns of original data while the remaining are black. Pyramid Adversarial Training [43] generates adversarial samples and perform a matched Dropout technique and stochastic depth regularization to enhance the robustness.
@@
The concept of perturbation through the addition of noise is also employed in machine unlearning tasks. [44] utilizes random Gaussian noise perturbations to the source model $M$’s parameters, constrained by ($\epsilon,\delta$), ensuring that the model remains indistinguishable from any point within a small neighborhood of the optimal model.
@@
Where $f_{m}$ denotes a random mechanism, with model parameters as its input. Additionally, $S$ represents the subset of the model parameter space. In the status quo, existing adversary-based unlearning methods concentrate on ($\epsilon,\delta$)-perturbations, which are based on perturbing all parameters of the model.
@@ III-A Problem Definition
In this paper, we focus on machine unlearning problem. Let $D$ represent the training data set which contains $n$ samples $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),...,\left(x_{n},y_{n}\right)$. For each sample, $x\in X$ and $y\in Y$. These data are posted by multiple users and collected by some institutions or organization and then labeled. Suppose $D=D_{UL}\cup D_{RE}$ and $D_{UL}\cap D_{RE}=\emptyset$, here $D_{UL}$ represents the unlearning data, which is absolutely randomly chosen by users to be unlearned and $D_{RE}$ represents the remaining data after unlearning data removal. Given a trained model $M$ with parameter $w$ on data set $D$, when conducting machine unlearning, our goal is to gain a target model $M_{UL}$ with parameter $\theta$ while the number of perturbed parameters should be as few as possible, and this $M_{UL}$ eliminates the influence of $D_{UL}$ and holds a quite good performance on $D_{RE}$. Suppose $w^{\prime}$ is the parameter after fine-grained perturbation. And as previously mentioned, we assume $P(M,D)$ as the performance of a model $M$ on data $D$, thus, the effect that our fine-grained perturbation machine unlearning aims to is :
@@
$P(M_{\theta},D_{RE})\approx P(M_{w},D_{RE})$
$minize\|w^{\prime}-w\|_{0}$ (3)
where $\|w^{\prime}-w\|_{0}$ is $L_{0}$ norm, representing the number of perturbed parameters.
Our fine-grained perturbation machine unlearning includes two stages: perturbation and fine-tuning. And the similarity of the output distributions of the model $M$ and $M_{UL}$ guides whether the fine-tuning should be terminated. We assume the posterior distribution of $M$ and $M_{UL}$ are $M\left(D_{UL}\right)\sim P_{1}$ and $M_{UL}\left(D_{UL}\right)\sim P_{2}$. When $P_{1}$ and $P_{2}$ are quite similar, we claim machine unlearning is conducted.
III-B Our Unlearning Approachs
The parameters in the machine learning model store the high-dimensional features of the data, so in order to eliminate the influence of unlearning data on the model, inspired by adversarial learning [36], we consider adding slight perturbation to trained model and then perform several fine-tuning epochs. Current machine unlearning methods based on parameters perturbation basically change all the parameters of the model or use tricks such as regularization or Dropout to improve the efficiency of unlearning. We desire to perturb the parameters of the model at a finer granularity.
@@ III-B Our Unlearning Approachs
Not all parameters affect the model equally. For instance, for images, the impact of the former layers on the model is mainly to extract the features (like color, shape, etc) of the data, while the later layers obtain high-dimensional information, thereby completing tasks such as object recognition. To this extent, it can be considered that perturbing the last few layers and freezing the former layers can achieve the purpose of inexact machine unlearning. However, only modifying the last few layers leads to a lower degree of machine unlearning, as shown in Tab.II, where the accuracy $Acc_{UL}$ on $D_{UL}$ for EU-5, EU-10, CF-5, CF-10 is significantly higher than other unlearning strategies.
In order to make the perturbation ratio decrease and make the perturbation more precise, we consider identifying several parameters in the model that have the greatest impact on performance (which we call ’most sensitive’) to perturb, or to address the issue of reduced unlearning effect caused by perturbing the parameters of the last few layers. Due to the dependencies among the model’s parameters, we adopt a fine-grained random perturbation unlearning strategy ignoring these dependencies, aiming to explore the differences among the most sensitive unlearning strategy and fine-grained random perturbation strategy. Thus, we propose two strategies for perturbing model parameters, namely Random-k and Top-K.
@@ III-B1 Parameter Selection and Perturbation
For Random-k, by setting the perturbation ratio $k$, randomly select the model parameters to be perturbed. For Top-K, we intend to calculate the parameter sensitivity on the data set to screen out the Top-K parameters that have the greatest impact on the model, so as to achieve the purpose of unlearning by performing parameter perturbation in a lighter way.
For the Top-K strategy, as previously mentioned, it is necessary to evaluate the sensitivity of each model parameter. Certainly, perturbing model parameters one by one and rank their performance change can determine the most sensitive $k$ parameters, but it’s time-consuming. To quickly obtain the sensitivity of parameters, we approximate it using the gradient. Specifically, after the model has converged, yielding the function $F(x,w)$, we introduce a small perturbation to the model parameters denoted as $\epsilon=\{\epsilon_{i,j}\}$. The change in the model’s output can be approximately measured using the following formula:
$\centering\bigtriangleup(x,w)=\sum_{i,j}\frac{\partial F(x,w)}{\partial w_{i,j}}*\epsilon_{i,j}\@add@centering$ (4)
Where $\bigtriangleup(x,w)$ represents the change in the model’s output performance due to the introduced perturbation $\epsilon$, and $\partial F(x,w)/\partial\theta_{i,j}=g_{i,j}(x)$ is the gradient of learned model to parameter $w$. This evaluation of sensitivity allows us to assess the impact of each parameter on the model’s predictions.
Focusing on aforementioned Equation 4, assuming the perturbation $\epsilon$ added to the model is a same constant for all parameters, the calculation of sensitivity can be simplified as follows:
@@
By adding such a constant perturbation across all model parameters, we can directly evaluate the sensitivity of each parameter’s contribution to the model’s overall output change. Only with one forward pass and back-propagation process, we obtain the sensitivity for each parameter.
@@
$w_{i}=w_{i}+\epsilon*w_{i}$
$w_{i}\in\{Top-K(w)\lor{Random-k(w)}\}$ (6)
Here $\epsilon$ is a coefficient with a small value. By introducing these perturbations, the model’s learned information is rendered inexact and less influential, allowing it to gradually fade out without abrupt removal. This approach offers a practical and robust solution for machine unlearning, enabling models to efficiently adapt to changing data while preserving essential knowledge from previous training.

### v4

@@ Abstract
However, commonly used training data are independent and identically distributed, for inexact machine unlearning, current metrics are inadequate in quantifying unlearning degree that occurs after unlearning. To address this quantification issue, we introduce SPD-GAN, which subtly perturbs data distribution targeted for unlearning. Then, we evaluate unlearning degree by measuring the performance difference of the models on the perturbed unlearning data before and after unlearning. Furthermore, to demonstrate efficacy, we tackle the challenge of evaluating machine unlearning by assessing model generalization across unlearning and remaining data. To better assess the unlearning effect and model generalization, we propose novel metrics, namely, the forgetting rate and memory retention rate. By implementing these innovative techniques and metrics, we achieve computationally efficacious privacy protection in machine learning applications without significant sacrifice of model performance. A by-product of our work is a novel method for evaluating and quantifying unlearning degree.
@@ I Introduction
Users may want the removal of personal data in possession of an organization, typically deleting it randomly based on their specific needs. In several jurisdictions, this has compliance implications under regulations such as the General Data Protection Regulation (GDPR) in the European Union [1] and the California Consumer Privacy Act in the United States [2]. Machine learning models trained using such data retain their influence, thus risking privacy compromises [3]. For instance, in a recommendation system using collaborative filtering [4, 5, 6], if the model does not eliminate the influence of removed data (which we call ‘unlearning data’), it may continue to make recommendations based on user similarity.
@@
Differently, inexact machine unlearning [12, 13] aims to reduce time and resources requirements by selectively eliminating the data’s influence on the model. Current inexact machine unlearning methods based on model weight perturbation typically involve adding random Gaussian noise to all model parameters and then training for several epochs based on remaining data. We propose to achieve inexact unlearning using a more fine-grained perturbation. To that end, we design two inexact machine unlearning strategies: Random-k and Top-K. Here, Random-k represents perturbing kof the parameters, while Top-K represents perturbing the top K parameters. To avoid notational confusion, we distinguish the two approaches by using lower and upper case respectively. By selectively perturbing a small subset of parameters, we aim to achieve the desired effect of machine unlearning efficiently while also reducing the impact on the overall model performance.
@@
However, the existing metrics are unable to reveal the extent of unlearning. Suppose splitting training data into two parts: unlearning data $D_{UL}$ and remaining data $D_{RE}$, where $D=D_{UL}\cup D_{RE}$, $|D_{UL}|\ll|D_{RE}|$. As shown in Fig.1, which shows an early glimpse of some results achieved by one of our own unlearning techniques), while the accuracy may vary across $D_{UL}$ and $D_{RE}$ for each model, it consistently maintains a high accuracy for $D_{UL}$. Consequently, it becomes challenging to quantify the extent to which the model eliminates the influence of $D_{UL}$. Since the performance difference of unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$ is relatively small. We attribute this to two reasons, first, a good generalization capability of the model, second, the characteristic that the training data followed an independent and identically distributed (i.i.d) property.
In order to amplify the performance difference between the unlearning model $M_{UL}$ on $D_{UL}$ and $D_{RE}$, we propose a SPD-GAN (Slightly Perturbed Distribution-Generative Adversarial Network) to break the i.i.d property of the unlearning data. For the unlearning data $D_{UL}$, we introduce perturbations to it’s distribution, thus we gain perturbed unlearning data $D_{p}$. We limit the perturbations to be very slight so that the accuracy of source model $M$ on $D_{p}$ is close to the accuracy of $M$ on $D_{UL}$. Suppose $P(M,D)$ represents the performance of model $M$ on data $D$, seeing in the Eq.1, if the perturbed unlearning data $D_{p}$ exhibits significant performance difference between source model and unlearning model, we can then obtain the unlearning degree of each unlearning strategies.
@@ II-A User Privacy
Federated learning (FL) is a decentralized approach to protect user privacy, where training data remains distributed across multiple devices and servers, and the model is trained without centralizing the data [28]. FL methods based on differential privacy have been proposed to protect user privacy, with the principal focus on attacks and defenses [29] in FL systems. Furthermore, recent advancements in FL incorporate various cryptographic techniques, such as secure multi-party computation and homomorphic encryption, to enhance the security and privacy of the training process. This ongoing research strengthens the privacy-preserving capabilities of FL.
@@ II-B Machine Unlearning
Several approaches have emerged to address the limitations of retraining models from scratch for machine unlearning, since it is generally impractical. the most common approach is SISA (Sharded, Isolated, Sliced, and Aggregated) training [8], which splits data into shards and trains them in isolation for speed and effectiveness when unlearning. Liu et al. [31] combine continual learning with private learning inspired by the SISA approach, but SISA-based methods often consume a lot of storage. Following the combination of machine unlearing and continual learning, Zuo et al. [32] propose a embedding-based framework, which modifies the data rather than the model itself to achieve unlearning goal by using vector databases.
@@
In addition to combining with continual learning, machine unlearning is also integrated with other approaches, such as federated learning, forming a research area known as federated unlearning [34] [35]. In federated unlearning, considerations are made for clients leaving the federation or for unlearning data corresponding to clients. For instance, Liu et al. [34] propose FedEraser, which unlearns a client’s data by reconstructing an unlearned global model instead of retraining from scratch. Zhao et al. [35] utilize knowledge distillation techniques to recover the contribution of client models to achieve federated unlearning.
@@ II-C Adversarial Learning
Moreover, these techniques have paved the way for numerous subsequent studies in adversarial machine learning, illustrating the continuous evolution of this field. In other words, leveraging adversarial training enhances the robustness of the model and better identify and defend against malicious attacks. Smoothed ViTs [44] are trained using only one columns of original data while the remaining are black. Pyramid Adversarial Training [45] generates adversarial samples and perform a matched Dropout technique and stochastic depth regularization to enhance the robustness.
The concept of perturbation through the addition of noise is also employed in machine unlearning tasks. Neel et al.[46] utilizes random Gaussian noise perturbations to the source model $M$’s parameters, constrained by ($\epsilon,\delta$), ensuring that the model remains indistinguishable from any point within a small neighborhood of the optimal model.
@@
Where $f_{m}$ denotes a random mechanism, with model parameters as its input. Additionally, $S$ represents the subset of the model parameter space. In the status quo, existing adversary-based unlearning methods concentrate on ($\epsilon,\delta$)-perturbations, which involve perturbing all the parameters of the model.
@@ III-A Problem Definition
In this paper, we focus on machine unlearning problem. Let $D$ represents the training data set which contains $n$ samples $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),...,\left(x_{n},y_{n}\right)$. For each sample, $x\in X$ and $y\in Y$. These data are posted by multiple users, collected by institutions or organizations, and then labeled. Suppose $D=D_{UL}\cup D_{RE}$ and $D_{UL}\cap D_{RE}=\emptyset$, here $D_{UL}$ represents the unlearning data, which is absolutely randomly chosen by users to be unlearned and $D_{RE}$ represents the remaining data after unlearning data removal.
Given a trained model $M$ with parameter $\omega$ on data set $D$, when conducting machine unlearning, our goal is to gain a target model $M_{UL}$ with parameter $\theta$ while the number of perturbed parameters should be as few as possible, and this $M_{UL}$ eliminates the influence of $D_{UL}$ and holds a quite good performance on $D_{RE}$. Suppose $\omega^{\prime}$ is the parameter after fine-grained perturbation. During fine-tuning, the parameters are represented by $\theta$. And as previously mentioned, we assume $P(M,D)$ as the performance of a model $M$ on data $D$, thus, the effect that our fine-grained perturbation machine unlearning aims to is :
@@
$P(M_{\theta},D_{RE})\approx P(M_{\omega},D_{RE})$
$minimize~{}\|\omega^{\prime}-\omega\|_{0}$ (3)
where $\|\omega^{\prime}-\omega\|_{0}$ is $L_{0}$ norm, representing the number of perturbed parameters.
We redefine the objective as a loss function, aiming to maximize the prediction error on unlearning data by guiding the model to make predictions more random or close to its untrained state. At the same time, it is necessary to maintain the performance on the remaining data as much as possible.
$\mathcal{L}=\mathbb{E}_{x\in D_{UL}}\|\nabla_{\theta}\ell(f(x),y)\|^{2}-\mathbb{E}_{x\in D_{RE}}\|\nabla_{\theta}\ell(f(x),y)\|^{2}$ (4)
Our fine-grained perturbation-based machine unlearning involves two stages: perturbing and fine-tuning. After applying perturbations to the model, the similarity of the output distributions of the model $M$ and $M_{UL}$ guides whether the fine-tuning should be terminated. Assuming the posterior distribution of $M$ and $M_{UL}$ are $M\left(D_{UL}\right)\sim P_{1}$ and $M_{UL}\left(D_{UL}\right)\sim P_{2}$. When $P_{1}$ and $P_{2}$ are quite similar, we claim machine unlearning is conducted.
III-B Our Unlearning Approaches
The parameters in the machine learning model store the high-dimensional features of the data, so in order to eliminate the influence of unlearning data on the model, inspired by adversarial learning [38], we consider adding slight perturbation to trained model and then perform several fine-tuning epochs. Current machine unlearning methods based on parameters perturbation basically change all the parameters of the model or use tricks such as regularization or Dropout to improve the efficiency of unlearning. In contrast, we aim to perturb the parameters of the model at a finer granularity.
@@ III-B Our Unlearning Approaches
Not all parameters affect the model equally. For instance, in image processing tasks, the earlier layers primarily focus on extracting features such as color and shape, while the later layers obtain high-dimensional information, thereby completing tasks such as object recognition. To this extent, it can be considered that perturbing the last few layers and freezing the former layers can achieve the purpose of inexact machine unlearning.
However, only modifying the last few layers leads to a lower degree of machine unlearning, as shown in Tab.II, where the accuracy $Acc_{UL}$ on $D_{UL}$ for EU-5, EU-10, CF-5, CF-10 is significantly higher than other unlearning strategies.
In order to make the perturbation ratio decrease and make the perturbation more precise, and to address the issue of reduced unlearning effect caused by perturbing the parameters of the last few layers. we consider identifying several parameters in the model that have the greatest impact on performance (which we call ’most sensitive’) to perturb. Thus, we propose Top-K unlearning strategy.
However, the Top-K unlearning strategy usually requires sensitivity calculations, while ignoring parameters with lower sensitivity that may still contribute to the unlearning effect. Moreover, the Top-K unlearning strategy overlooks the dependencies between parameters. Hence, we adopt a fine-grained random perturbation unlearning strategy (which we call ”Random-k”) ignoring these dependencies, aiming to explore the differences among the most sensitive unlearning strategy and fine-grained random perturbation strategy.
As shown in Fig.2, the proposed unlearning strategy consists of two processes: perturbing and fine-tuning. The details include two components: the selection of perturbed parameters and the unlearning process.
@@ III-B1 Parameter Selection and Perturbation
For Random-k, by setting the perturbation ratio $k$, randomly select the model parameters to be perturbed. For Top-K, we intend to calculate the parameter sensitivity on the data set to identify the Top-K parameters that have the greatest impact on the model, so as to achieve the purpose of unlearning by performing parameter perturbation in a lighter way.
For the Top-K strategy, as previously mentioned, it is necessary to evaluate the sensitivity of each model parameter. Certainly, perturbing model parameters one by one and rank their performance change can determine the most sensitive $k$ parameters, but it’s time-consuming. To quickly obtain the sensitivity of parameters, we approximate it using the gradient. A similar approach is also reflected in [47] for continual learning. Specifically, after the model has converged, yielding the function $F(x,w)$, we introduce a small perturbation to the model parameters denoted as $\mu=\{\mu_{i,j}\}$. The change in the model’s output can be approximately measured using the following formula:
$\centering\bigtriangleup(x,\omega)=\sum_{i,j}\frac{\partial F(x,\omega)}{\partial\omega_{i,j}}*\mu_{i,j}\@add@centering$ (5)
Where $\bigtriangleup(x,w)$ represents the change in the model’s output performance due to the introduced perturbation $\epsilon$, and $\partial F(x,w)/\partial\omega_{i,j}=g_{i,j}(x)$ is the gradient of learned model to parameter $w$. This evaluation of sensitivity allows us to assess the impact of each parameter on the model’s predictions.
Focusing on aforementioned Eq. 5, assuming the perturbation $\epsilon$ added to the model is a same constant for all parameters, the calculation of sensitivity can be simplified as follows:
@@
By adding such a constant perturbation across all model parameters, we can directly evaluate the sensitivity of each parameter’s contribution to the model’s overall output change, which can be approximated by calculating the norm of the gradient. Only with one forward pass and back-propagation process, we obtain the sensitivity for each parameter.
@@
$\omega_{i}=\omega_{i}+\epsilon*\omega_{i}$
$\omega_{i}\in\{Top-K(\omega)\lor{Random-k(\omega)}\}$ (7)
Here $\epsilon$ is a coefficient with a small value. By introducing these perturbations, the model’s learned information is rendered inexact and less influential, allowing it to gradually diminish rather than being abruptly removed. This approach offers a practical and robust solution for machine unlearning, enabling models to efficiently adapt to changing data while preserving essential knowledge from previous training.